Data concerning Wislocki Corporation's single product appear below:

Per Unit Percent of Sales
Selling price $ 190 100 %
Variable expenses 38 20 %
Contribution margin $ 152 80 %

Fixed expenses are $1,039,000 per month. The company is currently selling 9,800 units per month.

Required:

The marketing manager would like to introduce sales commissions as an incentive for the sales staff. The marketing manager has proposed a commission of $13 per unit. In exchange, the sales staff would accept an overall decrease in their salaries of $107,000 per month. The marketing manager predicts that introducing this sales incentive would increase monthly sales by 360 units. What should be the overall effect on the company's monthly net operating income of this change?
Answer #1

Total Contribution margin - Fixed cost = Net operating income

Present net operating income

= (9800*152) - 1,039,000 = 450,600

Proposed net operating income

= (10,160*(190-51)) - (1,039,000-107,000)

= 480,240

Increase in net operating income = 480,240 - 450,600

Increase in net operating income = 29,640